k, 6.124 1, 2S8 6693 6Ilf 2. 298 (ii) A first order reaction is 75.0% complete inamo calculate the. half- lives for this reaction and also calculate the time required to complete 90.0% of the reaction.


PLEASE SHOW ALL STEPS CLEARLY
Answer #1

ln A = -kt + ln Ao

k = rate constant

t = time

Ao = Initial concentration

A = concentration at time t

Let us find rate constant using the values

ln 0.25 = - k 320 + ln 1

k = 0.004332 sec-1

Now let us calculate the haf life

half life = 0.693 / k

First Half Life = 0.693 / 0.004332   = 159.96 Seconds

Second Half Life = 2 x 770.85 = 319.93 Seconds

time to complete 90% reaction

ln 0.1 = -0.004332 x t + ln 1

t = 531.5 seconds

It will take 531.5 seconds to complete 90% of the reaction